OpenStudy (anonymous):

If sinθ = 3/4 what are the values of cos θ

OpenStudy (anonymous):

Do you know the Pythagorean theorem?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

you can use that to find your adjacent side. Once you find that, you can find your cos

OpenStudy (anonymous):

so \[3^{2}+b ^{2}=4^{2} \]to find the adjacent side

OpenStudy (anonymous):

Okay, go on

OpenStudy (anonymous):

\[b=\sqrt{5}\] so what would cos be?

OpenStudy (anonymous):

2.23

OpenStudy (anonymous):



OpenStudy (anonymous):

right. I would leave it as sqrt of 5 unless it tells you to find the sqrt

OpenStudy (anonymous):

Oh okay, that's the answer then?

OpenStudy (anonymous):

well, it's asking for cos. so from that picture, what is cos of theta?

OpenStudy (anonymous):

I don't know how to find that..

OpenStudy (anonymous):

\[\cos \theta=\frac{ adj }{ hyp }\]\[\cos \theta=\frac{ \sqrt{5} }{ 4 }\]That's your answer

OpenStudy (anonymous):

Awesome! Is that it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you!